OpenStudy (anonymous):

Need algebra 2 help!!!

OpenStudy (anonymous):

If x>0, the expression (√x)(√2x) is equivalent to,,.

geerky42 (geerky42):

\(\sqrt{x}\sqrt{2x} = \sqrt{(x)(2x)}\), right?

OpenStudy (austindeboe18):

geerky is right